# FakeScope: Large Multimodal Expert Model for Transparent AI-Generated Image Forensics

### v1

@@ VI-A3 Baseline Models
The proposed FakeScope is compared with a diverse set of models, including open-sourced LMMs (InstructBLIP [76], IDEFICS-Instruct [77], InternLM-XComposer2-vl [31], LLaVA-v1.5 [36], Qwen-VL [78], mPLUG-Owl2 [37]), proprietary LMMs (GPT-4V [33], GeminiPro [34], Claude3-Sonnet [35]), data-driven binary forensic models (CNNSpot [20], UnivFD [3], FreDect [22]), and the vision-language contrastive model CLIP-ViT-Large-14 [79]. For detection evaluation, LMMs supports both qualitative (textual judgments) and quantitative (probabilistic estimation) settings, whereas binary detectors are limited to the quantitative mode. To ensure fair comparisons, the data-driven models are retrained on the same images as FakeScope without data augmentation, and all tested LMMs are evaluated using same prompting templates as FakeScope. These settings remain consistent across all experiments unless explicitly stated otherwise.
TABLE III: Forensic detection results on the FakeBench-FakeClass [7] dataset under the qualitative setting. The detection Acc (%) of each model is reported. The first and second best LMMs are highlighted in bold and underlined, respectively. The “Fake” column reports the average accuracy across all individual “Generation Model” columns, and both the averaged “Authenticity” columns and “Question Type” columns reflect the “Overall” column.
@@ VI-A4 Performance Measures
FakeScope is versatile in coping with forensic detection and in-depth analysis simultaneously. To comprehensively evaluate its multidimensional capabilities, we adopt a range of metrics. For forensic detection, we consider both qualitative and quantitative settings. Specifically, qualitative classification accuracy is measured by $Acc=N_{c}/N_{t}$, where $N_{c}$ and $N_{t}$ denote the number of correct and total predictions, respectively. For quantitative assessment, we follow [20, 3] and use threshold-independent calibration metrics, including average precision (AP) and area under the receiver operating characteristic curve (AUC), which are robust to class imbalance across datasets. To assess transparency-related performance, we evaluate the alignment between model responses and reference answers by employing both conventional word-level metrics (i.e., BLEU [80], ROUGE [81], and BERTScore [82]) and the LLM-as-a-judge strategy [83] to measure semantic-level alignment, following [40, 36, 7, 16]. The measures are then normalized and averaged to obtain the overall reflection, where higher values indicate better performance on transparency-related abilities. Details on evaluation measures are provided in SM.
Figure 6: Forensic detection results on FakeClass [7] under the quantitative setting. Calibration measures Acc (threshold is set to 0.5 for balanced indication), AP, and AUC (%) are reported for each model. The FakeScope far exceeds the other models on each subset and the entire dataset.
@@ VI-B Multimodal Forensic Capability
The multimodal forensic capabilities of the proposed FakeScope model are quantitatively assessed on FakeBench containing FakeClass, FakeClue, and FakeQA subsets with respect to four ability criteria for LMMs comprehensively defined in [7], which are stated as follows.
@@
  • Fine-grained Analyzing, which investigates whether LMMs can discuss and analyze specific aspects of telltale visual clues.
@@
TABLE IV: Multimodal forensic capability comparisons on FakeBench [7], including detecting on FakeClass, interpreting on FakeClue-faultfinding, reasoning on FakeClue-inference, and fine-grained forgery analyzing on FakeQA. The B.-1/B.-2 denote BLEU-1/2, R.-L is ROUGE-L, Sim. is BERTScore similarity, and $\alpha$, $\rho$, $\kappa$ correspond to completeness, preciseness, and relevance judged by LLMs. The first- and second-best performances of each column are highlighted in bold and underlined, respectively.
@@
LMMs Word-level Semantic-level Avr.$\uparrow$ Word-level Semantic-level Avr.$\uparrow$ Word-level Semantic-level Avr.$\uparrow$
@@
1) Detection Task: Simple detection of fake and real images is the fundamental capability of forensic detectors. In Table III, we compare the performance of the proposed FakeScope with other LMMs under the qualitative setting on the FakeClass dataset. The FakeClass dataset contains 6,000 diversified query-answer pairs for judging absolute image authenticity with balanced authenticity proportions and adequate query diversity. The detection accuracy is measured by the correctness of responses to these queries. As reflected in the results, FakeScope of the two versions simultaneously exhibits noticeable improvements in detection reliability across all generators, authenticity categories, and query types, suggesting that FakeInstruct can significantly enhance the forensic awareness of LMMs on binary classification. Notably, FakeScope achieves a peak accuracy of 99.68%, outperforming even the best-performing human expert, and marking a significant milestone in automated forensic analysis. In Fig. 6, we further present the performance comparison using calibration measures under the quantitative setting. In this case, one can still witness a consistently significant advantage across all categories over both other LMMs, indicating the effectiveness of the proposed probability estimation strategy based on token soft-scoring. Notably, since we do not explicitly incorporate numerical values as the supervision when training LMMs as forensic detectors, the results further suggest that LMMs can effectively adapt to hard authenticity labels via the internalized probabilistic reasoning process and perform well on estimating numerical probabilities. In particular, despite being trained with the same image data, CNNSpot [20], UnivFD [3], and FreDect [22] exhibit inferior performance compared with FakeScope. While these data-driven methods can be effective in some scenarios, their performance often benefits from elaborate data augmentation [20]. In contrast, under the controlled setting of this study where no image-level augmentation is applied, purely training with binary labels may fall short in capturing the intricacies of forensic cues. This further underscores the effectiveness of the multimodal information in FakeInstruct and highlights LMMs’ superiority to integrate multimodal information for AI-generated image forensics.
@@
2) Interpreting Task: We find out that incorporating visual trace evidence into training can significantly enhance the interpretation ability of LMMs. In practice, LMMs are evaluated using the fault-finding prompting mode of FakeClue, where models are instructed to explain image authenticity based on visual clues, such as in the prompt “This is a fake image generated by AI, explain the reasons”. In the first category of Table IV, the two variants of FakeScope demonstrate substantial improvements in interpreting ability, especially the word-level similarities and semantic relevance level to the ground truth. These findings indicate that training on FakeInstruct conspicuously enhances the effect-to-cause awareness concerning the forensic attributes of LMMs.
@@
3) Reasoning Task: Forensic reasoning is the inverse of interpretation, assessing LMMs’ inference capability in deriving authenticity judgments from visual trace evidence. This task is evaluated using the prompts in the inference mode of FakeClue, such as “Inspect this image carefully and formulate an evaluation regarding its authenticity”. As observed in the experimental results listed in the second category of Table IV, while all baseline LMMs generally exhibit weak forensic reasoning capabilities, the proposed FakeScope achieves a substantial improvement after training on FakeInstruct and even surpasses all compared models, reaching SOTA performance. Most notably, its reasoning capability also outperforms the teacher model GPT-4V [33] by over 30%, demonstrating the effectiveness of FakeInstruct in enhancing cause-to-effect forensic awareness and triggering forensic reasoning in LMMs.
4) Fine-Grained Analyzing Task: To further evaluate the effectiveness of FakeScope in analyzing specific aspects of image forgery, we conduct experiments on the FakeQA dataset, which consists of queries regarding fine-grained visual forgeries. As shown in the third category of Table IV, FakeScope exhibits notable improvements over both baseline models and its teacher model, highlighting its enhanced fine-grained forensic analyzing capabilities. However, despite these promising results, this ability still shows room for improvement compared to the other two transparency-related forensic capabilities interpreting and reasoning. This suggests potential avenues for further optimization, particularly in refining its sensitivity to subtle forgery artifacts.
TABLE V: In-the-wild forensic detection ability on WildRF [73] and SynthWildX [4] containing user-generated image content and unknown generators. The AP and AUC (%) are reported, and the first- and second-best performers for each category are highlighted in bold and underline respectively.
@@
TABLE VI: Generalization capability validated on AntifakePrompt [71] and AIGCDetectBenchmark [72]. The AP (%) of various classifiers across 17 individual generators are reported. Chance is 50% and best possible performance is 100%. All the LMMs are tested under the quantitative setting. Values in gray represent the generators seen in training. No content overlapping exists in the results. The best model in each column is highlighted in bold.
@@
CNNSpot 64.86 61.41 70.33 60.10 63.08 56.41 70.19 82.85 92.51 86.50 56.61 60.25 48.88 52.15 50.54 50.22 95.15 60.71 66.00
UnivFD 61.37 50.36 39.78 55.61 48.49 51.91 56.22 43.25 35.75 44.24 45.36 58.17 50.85 57.46 80.99 60.68 68.42 57.24 53.47
FreDect 49.93 55.13 49.99 53.26 54.93 49.97 53.61 29.79 37.78 38.36 35.80 44.46 49.86 49.93 48.02 49.81 56.10 50.63 47.45
CLIP-ViT-Large-14 89.22 71.12 94.92 88.58 64.40 81.17 60.43 29.20 31.87 37.28 49.44 70.85 62.86 63.26 85.15 71.71 41.85 71.38 64.31
InternLM-XC.2-vl 52.15 46.40 61.16 50.47 52.28 44.29 68.72 35.73 41.03 48.59 45.49 62.15 52.39 58.33 60.58 58.46 71.70 58.23 53.52
InstructBLIP 50.27 40.79 32.31 88.63 40.42 49.36 55.46 25.98 29.87 25.86 72.11 64.04 59.18 74.46 95.64 81.84 97.48 67.17 57.86
Qwen-VL 72.27 42.91 41.39 82.54 44.80 48.92 84.18 75.39 69.55 72.97 83.83 88.15 66.50 78.31 93.55 89.05 91.31 73.52 72.10
IDEFICS-Instruct 40.26 51.87 53.68 39.40 47.26 66.99 35.16 22.03 23.55 25.85 22.45 41.24 46.53 41.11 34.81 44.27 48.49 45.36 40.29
mPLUG-Owl2 58.44 48.91 83.24 51.39 45.26 41.84 69.79 45.10 46.10 44.71 55.03 74.34 50.71 57.82 77.64 68.28 71.16 62.86 58.22
LLaVA-v1.5 40.97 44.09 41.30 38.08 43.03 56.01 42.36 24.32 25.07 26.71 32.78 45.13 47.93 45.08 37.69 43.91 80.87 47.40 42.08
@@ VI-C In-the-wild Performance
To evaluate the real-world applicability of our model, we conduct in-the-wild experiments on diverse, unstructured image data collected from social media, including WildRF [73] and SynthWildX [4]. These datasets contains real-world samples with various distortions, such as compression artifacts, post-processing modifications, and noise, which pose additional challenges for forensic models. The testing protocol aligns with other experiments, and all LMMs are evaluated in the quantitative setting. The AP and AUC values for detection are listed in Table V. As reflected by the results, FakeScope equipped with mPLUG-Owl2 [37]consistently outperforms other state-of-the-art methods across the majority of generators, achieving superior overall performance in the in-the-wild testing. Notably, it demonstrates competitive results on images generated by Dalle3 and Midjourney, ranking second only to CNNSpot [20]. These findings highlight FakeScope’s robustness and effectiveness in handling unconstrained image content and generation models, reinforcing its adaptability to real-world forensic challenges beyond in-lab benchmarks.
Figure 7: Accuracy as a function of detection threshold. For each model, we present the relations between the detection threshold (x-axis) and accuracy (y-axis) on all generators of five testing databases. It turns out the FakeScope of two versions of base models (the last two sub-figures) is the only one that can achieve a great bargain between threshold sensitivity and detection accuracy, indicating remarkable robustness.
Figure 8: Performance on FakeBench [7] with respect to training data scale, demonstrating the effectiveness of FakeInstruct data. The x-axis indicates the proportion of FakeInstruct used for training and the y-axis is the corresponding performance indicator of each ability dimension.
@@ VI-D Generalization Capability
For forensic detectors, the generalization ability on unacquainted image contents and generators is crucial for practical deployment. To assess this, we conduct a more challenging experiment to evaluate whether FakeScope is well generalizable on unseen image content and generators. Practically, we test all methods on the uncontaminated content proportions of AntifakePrompt [71] and AIGCDetectBenchmark [72], in comparison with the data employed in FakeInstruct. In particular, the AntifakePrompt dataset covers several OOD tasks beyond the primary scope of this work, such as image super-resolution, face spoofing [75], and inpainting, which pose a greater challenge to model generalization and robustness. In Table VI, the AP and AUC scores of each model across 17 image generators are presented, including 6 seen and 11 unseen ones. We also present the average AP over unseen generators as well as over all generators. From Table VI, it can be observed that the two variants of the FakeScope can achieve SOTA performance across all 17 generators and, notably, outperforms all baselines on the set of unseen generators. These results demonstrate that FakeScope can effectively handle novel image content and generation techniques, indicating strong generalization capability.
Furthermore, we analyze model stability by plotting detection accuracy as a function of the decision threshold ($\in(0,1)$) for each method, thereby assessing their sensitivity to threshold selection across different generators. In Fig. 7, a separate curve is shown for each forensic detector by averaging detecting accuracy across all generators in each dataset. Unlike AUC and mAP, the accuracy (Acc) is highly sensitive to threshold variation, and models with unstable performance may suffer under suboptimal thresholds, leading to less robustness. It can be observed from Fig. 7 that both versions of FakeScope exhibit highly consistent peak accuracy across different datasets and maintain stable performance over a broad range of threshold values. In contrast, other models show considerable variability in optimal accuracy across datasets and are overly sensitive to specific threshold settings. This indicates that FakeScope possesses prominent robustness towards forensic threshold across different datasets and generators, thereby enhancing its applicability in real-world scenarios.

### v2

@@ VI-A3 Baseline Models
FakeScope is compared against a broad set of baselines: six open-source LMMs, including InstructBLIP [14], IDEFICS-Instruct [39], InternLM-XComposer2-VL [20], LLaVA-v1.5 [45], Qwen-VL [4], and mPLUG-Owl2 [78]; three proprietary LMMs, namely GPT-4V [54], GeminiPro [67], and Claude3-Sonnet [3]; three data-driven binary detectors, CNNSpot [68], UnivFD [53], and FreDect [24]; and the vision-language contrastive model CLIP-ViT-Large-14 [57]. For detection, LMMs are evaluated in both qualitative (textual judgment) and quantitative (probability estimation) settings, while binary detectors are evaluated only in the quantitative setting. To ensure fairness, the data-driven detectors are retrained on the same images as FakeScope. All LMMs are evaluated using the same prompting templates as FakeScope. These settings remain consistent across all the experiments, unless explicitly stated otherwise.
TABLE II: Forensic detection results on the FakeBench-FakeClass [40] under the qualitative setting. The detection Acc. (%) of each model is reported. The first and second best LMMs are highlighted in bold and underlined, respectively. The “Fake” column reports the average accuracy across all individual “Generation Model” columns, and both the averaged “Authenticity” columns and “Question Type” columns reflect the “Overall” column.
@@ VI-A4 Performance Measures
FakeScope is designed to handle both forensic detection and open-ended analysis, so we evaluate it with a diverse set of metrics. For forensic detection, we consider both qualitative and quantitative settings. In the qualitative setting, classification accuracy is defined as $Acc.=N_{c}/N_{t}$, where $N_{c}$ and $N_{t}$ are the numbers of correct and total responses. In the quantitative setting, following [68, 53], we use threshold-independent calibration metrics, including average precision (AP) and area under the ROC curve (AUC), which are more robust to dataset class imbalance. To assess transparency-related performance, we measure how well model responses align with reference answers both conventional word-level metrics (BLEU [55], ROUGE [10], and BERTScore [59]) and the LLM-as-a-judge strategy [26] to measure semantic-level alignment, following [50, 45, 40, 71]. These scores are normalized and averaged into an overall transparency score, where higher values indicate better performance. Forensic attribution is evaluated by accuracy on FakeBenchMCQ. Details on the evaluation measures are further provided in the SM.
@@ VI-B Multimodal Forensic Capability
Figure 6: Forensic detection results on FakeClass [40] under the quantitative setting. AP and AUC (%) are reported for each image generation model. The FakeScope far exceeds the other models on each subset and the entire dataset.
The multimodal forensic capabilities of the proposed FakeScope model are quantitatively assessed on FakeBench containing FakeClass, FakeClue, and FakeQA subsets concerning four ability criteria for LMMs comprehensively defined in [40], which are stated as follows.
@@
  • Fine-grained Analyzing, which investigates whether LMMs can discuss and analyze specific aspects of telltale visual clues;
  • Forensic Attributing, which investigate whether LMMs can spot specific forensic visual cues in images.
@@
TABLE III: Multimodal forensic capability comparisons on FakeBench [40], including detecting on FakeClass, interpreting on FakeClue-faultfinding, reasoning on FakeClue-inference, and fine-grained forgery analyzing on FakeQA. The -M and -L denote base model mPLUG-Owl2 and LLaVA-1.5, respectively. The B.-1/B.-2 denote BLEU-1/2, R.-L is ROUGE-L, Sim. is BERTScore similarity, and $\alpha$, $\rho$, $\kappa$ correspond to completeness, preciseness, and relevance judged by LLMs. Each column’s first- and second-best performances are highlighted in bold and underlined.
@@
[…]
@@
TABLE IV: Forensic attribution capability (%) on FakeBenchMCQ. The -M and -L denote base model mPLUG-Owl2 and LLaVA-1.5.
Model Question Type Attribution Level Overall
Yes/No What Where How Low Medium High
Random guess 50.13 37.60 24.27 26.93 34.50 37.80 31.90 34.73
GPT-4V (Teacher) 79.47 68.80 55.87 62.40 70.70 61.50 67.60 66.60
GeminiPro 70.37 66.20 52.13 70.80 65.20 62.60 66.80 64.88
Claude3 Sonnet 68.27 67.73 66.80 65.20 68.30 68.80 63.90 67.00
InstructBLIP 56.27 56.00 54.93 53.60 57.70 58.40 49.50 55.20
IDEFICS-Instruct 48.67 49.60 52.27 53.47 50.20 49.70 53.10 51.00
InternLM-XC.2-vl 69.73 68.13 66.27 59.47 66.30 65.50 65.90 65.90
LLaVA-v1.5 60.40 62.27 63.07 61.60 63.80 62.60 59.11 61.83
Qwen-VL 64.13 64.93 60.93 60.67 62.30 61.80 63.90 62.67
mPLUG-Owl2 70.53 66.93 62.27 64.13 64.80 66.30 66.80 65.97
FakeScope-M 75.47 79.73 78.53 80.27 76.80 80.50 78.20 78.50
Improvement +4.94 +12.80 +16.26 +16.14 +12.00 +14.20 +11.40 +12.53
FakeScope-L 79.47 78.00 78.27 75.20 82.80 73.70 76.70 77.74
Improvement +19.07 +15.73 +15.20 +13.60 +19.00 +11.10 +17.59 +15.91
1) Detection Task: Distinguishing real from fake images is the core capability of forensic detectors. We compare FakeScope with other LMMs under the qualitative setting using the FakeClass dataset, which includes 6,000 diverse QA pairs with balanced authenticity labels and varied queries. As shown in Table II, both versions of FakeScope consistently outperform other models across all subcategories, indicating that FakeInstruct notably enhances LMMs’ forensic detection capabilities. FakeScope reaches a peak accuracy of 99.68%, exceeding even the strongest human expert. Figure 6 further compares models under the quantitative setting using calibration-based metrics. FakeScope remains the strongest model, demonstrating the effectiveness of the token-based soft-scoring strategy. Notably, despite receiving no explicit numerical supervision during training, it aligns closely with hard labels and produces reliable probability estimates, suggesting strong internalized reasoning. Even when trained on the same image data, conventional detectors perform worse than FakeScope. Without image-level augmentation, these data-driven detectors degrade markedly [68], underscoring the value of FakeInstruct’s multimodal supervision and the stronger forensic integration ability of LMMs.
@@
2) Interpreting Task: We evaluate LMMs with the fault-finding prompting mode of FakeClue, where models are instructed to explain image authenticity based on visual descriptions, such as in the prompt “This is a fake image generated by AI, explain the reasons”. In the first category of Table III, FakeScope achieve clear gains in this task, which indicates that training on FakeInstruct conspicuously enhances the effect-to-cause awareness concerning the forensic attributes of LMMs.
@@
3) Reasoning Task: The forensic reasoning task is the inverse of interpretation, assessing LMMs’ inference capability in deriving authenticity judgments from visual trace evidence. This task is evaluated using the prompts in the inference mode of FakeClue, such as “Inspect this image carefully and formulate an evaluation regarding its authenticity”. As shown in the second category of Table III, baselines generally perform poorly on this task, whereas FakeScope improves substantially after training on FakeInstruct and achieves state-of-the-art results. Most notably, it surpasses the teacher model GPT-4V by more than 30%, highlighting FakeInstruct’s effectiveness in strengthening cause-to-effect forensic awareness.
4) Fine-Grained Analyzing Task: We assess FakeScope on FakeQA, which targets fine-grained visual forgery queries. As shown in the third category of Table III, FakeScope clearly outperforms both baseline and teacher models, demonstrating stronger fine-grained forensic analysis. However, this capability still lags behind its interpreting and reasoning performance, suggesting the future potential to improve its answering open-ended questions on subtle forgery artifacts.
Figure 7: Accuracy as a function of detection threshold. For each model, we present the relations between the detection threshold (x-axis) and accuracy (y-axis) on all generators of the five testing databases. The FakeScope of two versions of base models (the last two sub-figures) is the only one that can achieve a great bargain between threshold sensitivity and detection accuracy, indicating remarkable robustness.
TABLE V: In-the-wild performance comparison. The AP and AUC (%) are reported, and top two performers for each category are highlighted in bold and underline respectively.
@@
TABLE VI: Generalization capability comparison. The AP (%) of various classifiers across 17 individual generators is reported. Values in gray represent the generators seen in training. The top performer is highlighted in bold.
@@
CNNSpot [68] 64.86 61.41 70.33 60.10 63.08 56.41 70.19 82.85 92.51 86.50 56.61 60.25 48.88 52.15 50.54 50.22 95.15 60.71 66.00
UnivFD [53] 61.37 50.36 39.78 55.61 48.49 51.91 56.22 43.25 35.75 44.24 45.36 58.17 50.85 57.46 80.99 60.68 68.42 57.24 53.47
FreDect [24] 49.93 55.13 49.99 53.26 54.93 49.97 53.61 29.79 37.78 38.36 35.80 44.46 49.86 49.93 48.02 49.81 56.10 50.63 47.45
CLIP-ViT-Large-14 [57] 89.22 71.12 94.92 88.58 64.40 81.17 60.43 29.20 31.87 37.28 49.44 70.85 62.86 63.26 85.15 71.71 41.85 71.38 64.31
InternLM-XC.2-vl [20] 52.15 46.40 61.16 50.47 52.28 44.29 68.72 35.73 41.03 48.59 45.49 62.15 52.39 58.33 60.58 58.46 71.70 58.23 53.52
InstructBLIP [14] 50.27 40.79 32.31 88.63 40.42 49.36 55.46 25.98 29.87 25.86 72.11 64.04 59.18 74.46 95.64 81.84 97.48 67.17 57.86
Qwen-VL [4] 72.27 42.91 41.39 82.54 44.80 48.92 84.18 75.39 69.55 72.97 83.83 88.15 66.50 78.31 93.55 89.05 91.31 73.52 72.10
IDEFICS-Instruct [39] 40.26 51.87 53.68 39.40 47.26 66.99 35.16 22.03 23.55 25.85 22.45 41.24 46.53 41.11 34.81 44.27 48.49 45.36 40.29
mPLUG-Owl2 [78] 58.44 48.91 83.24 51.39 45.26 41.84 69.79 45.10 46.10 44.71 55.03 74.34 50.71 57.82 77.64 68.28 71.16 62.86 58.22
LLaVA-v1.5 [45] 40.97 44.09 41.30 38.08 43.03 56.01 42.36 24.32 25.07 26.71 32.78 45.13 47.93 45.08 37.69 43.91 80.87 47.40 42.08
@@
5) Forensic Attribution Task: We evaluate forensic attribution on FakeBenchMCQ, containing close-ended questions on specific visual evidence. As listed in Table IV, FakeScope achieves clear improvements over its base models, indicating that it can not only justify detection with long-form rationales but also identify concrete forensic evidence effectively. This further confirms that FakeInstruct successfully imparts visual forensic awareness to LMMs. Among all dimensions, the largest gain appears on where questions. We also observe larger gains on low- and medium-level attributes, suggesting the model could potentially be further improved by incorporating more conceptually high-level forensic knowledge.
@@ VI-C In-the-wild Performance
To assess real-world applicability, we evaluate on in-the-wild datasets built from social-media imagery, including WildRF [7] and SynthWildX [12]. These contain real-world degradations which pose additional challenges for forensic models. As reflected by the results in Table V, FakeScope equipped with mPLUG-Owl2 [78] consistently outperforms prior state-of-the-art methods across most generators and achieves the best overall in-the-wild performance. It is also competitive on Dalle3 and Midjourney, ranking second only to CNNSpot [68]. These findings highlight FakeScope’s robustness in handling real-world forensic challenges beyond in-lab benchmarks.
@@ VI-D Generalization Capability
The generalization to unacquainted image contents and generators is crucial for practical deployment. To assess this, we test all methods on the uncontaminated content proportions of AntifakePrompt [8] and AIGCDetectBenchmark [82], whose content is disjoint from the data used in FakeInstruct. AntifakePrompt further includes OOD tasks beyond our primary scope, such as image super-resolution, face spoofing [62], and inpainting, which stress both robustness and generalization. Table VI presents results across 17 image generators, including six unseen and 11 unseen ones. Both FakeScope variants achieve superiority across all 17 generators and outperform all baselines on the unseen subset, indicating strong generalization to novel content and generation techniques.
We further evaluate stability by plotting detection accuracy as a function of the decision threshold ($\in(0,1)$), measuring sensitivity to threshold selection across generators. In Fig. 7, each curve is obtained by averaging accuracy over all generators in each dataset. Unlike AUC and mAP, accuracy is highly threshold-dependent; unstable detectors can degrade substantially under suboptimal thresholds. As shown in Fig. 7 both FakeScope variants exhibit consistent peak accuracy and remain stable over a wide threshold range. In contrast, competing methods show larger shifts in their optimal thresholds and greater performance volatility. This suggests FakeScope is less sensitive to threshold choice across datasets and generators, improving robustness in real-world settings.